OpenStudy (anonymous):

solve for x on the interval 0 is less than or equal to x which is less than or equal to 2pi : sin(2x) = - square root of 3 divided by 2

OpenStudy (anonymous):

\[\sin(2x) = - \frac{ \sqrt{3} }{ 2 }\]

OpenStudy (anonymous):

\[x=\frac{ \sin^{-1} (-\frac{ \sqrt{3} }{ 2 }) }{ 2 }\]